OpenStudy (anonymous):

Part A: Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

Y=A+BX is a standard form, get equations in this form

OpenStudy (anonymous):

Wait what I Don't understand this at all @aryandecoolest

OpenStudy (anonymous):

see let A= fixed cost and B =variable cost, then you will get two equations after substitution of values... Find value of A and B

OpenStudy (anonymous):

\[285 = A +3B\] \[510 = A+ 6B\] solve for A and B

OpenStudy (anonymous):

how would I do that? Wouldn't it be like 285 = a + 3b -3 on both sides ect.. @aryandecoolest

OpenStudy (anonymous):

yeah!! solve for a and b by elimination method and then substitute back in basic equation y = a +bx.

OpenStudy (anonymous):

Sorry im one of those people that is scared of being wrong lol. Anyways so we would be at 282 - a + b right? @aryandecoolest

OpenStudy (anonymous):

see... you will get value of b=75, solve for a now

OpenStudy (anonymous):

Wait howd you get that? and 510 = a + 6b - 6 on both sides 504 = a + b ? @aryandecoolest

OpenStudy (anonymous):

Aah! no you can't do like that that's wrong....z

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)

OpenStudy (anonymous):

Kms I have no idea what im doing. @aryandecoolest

OpenStudy (anonymous):

you can't do like that...subtracting both sides cannot help to eliminate a variable term

OpenStudy (anonymous):

ok... see you will get b=75 and a=60

OpenStudy (mathmale):

@aryandecoolest : Please speak in terms of what coolcatsclan CAN and should do, not in terms of what he / she can't do. Thank you.

OpenStudy (anonymous):

substitute back now...

OpenStudy (mathmale):

that's better.

OpenStudy (anonymous):

oh! )

OpenStudy (yanasidlinskiy):

This is for Part A: The way he got 75 is like this: Let (3,285) and (6,510) represent the coordinate points, where x = number of days coat is rented for; and y = price of rental Given these 2 points, we could write the equation of the straight line. We need the slope. Slope = (510-285)/(6-3) = 225/3 = 75 Do you understand now?

OpenStudy (anonymous):

kindof :\ @YanaSidlinskiy

OpenStudy (anonymous):

and now like i said substitute in y= a+bx .... Replace a and b with their respective values in it

OpenStudy (yanasidlinskiy):

What do you not understand?

OpenStudy (anonymous):

so now you have to graph line.... y=75x+60

OpenStudy (anonymous):

sorry im replying slow, im trying to think (Its not working out too well) All of it @YanaSidlinskiy How would I do that @aryandecoolest

OpenStudy (mathmale):

@coolcatsclan : Please follow Yana's advice; it's right on target. First, calculate the slope of the straight line in question. Graphing this line, while not essential, would probably help you visualize what's happening here. The question asks you to determine the value of y (the total cost of renting the coat as a function of x, the number of days it is rented. Use the point-slope formula for the equation of a straight line to obtain the necessary equation. Finally, simplify this equation into slope-intercept form.

OpenStudy (yanasidlinskiy):

It's Ok:) Do you see how I got the numbers though?

OpenStudy (anonymous):

find y -intercept...which is 60 (y = mx+b, b = y-intercept) Take another point, say x = 1, y = 75(1) + 60 = 135